# Moduli spaces of twisted equivariant $G$-bundles over a curve

### v1

@@ 6.3. The moduli space is a projective variety
By the proof of [20, Proposition 2.3.5.17], $\sigma_{C^{*}}$ extends to a family of homomorphisms $\sigma_{C}:(E_{C})_{a,b,c}\to\pi_{X}^{*}\mathcal{O}_{X}$. As in such proof, let $\iota:(C\times X)\setminus(C\times Z)\hookrightarrow C\times X$ be the open embedding and let
@@
Finally, as in the proof of [20, Proposition 2.3.5.17], one can check that $(\widetilde{E}_{0},\widetilde{\sigma}_{0},\widetilde{f}_{0})$ is a $(\delta,\chi)$-semistable pseudo-equivariant pair. Since its image in $\mathbb{G}{(m)}^{ss}$ is equal to $\eta(0)$, it determines a extension $\eta^{*}:C\to\mathfrak{C}_{0}^{ss}$ lifting $\eta$, as required. ∎
@@ 6.4. S-equivalence vs polystability
We claim that $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)=0$ implies that $M(E_{t\bullet},\alpha_{\bullet})+\delta\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=0$. This is just the same calculation as for the moduli space of principal $G$-bundles with no twisted equivariant structure. For example, use a similar calculation to the one in the proof of [20, Theorem 2.3.5.13] after the statement of [20, Corollary 2.3.5.14].
@@
Now assume that $M(E_{t\bullet},\alpha_{\bullet})=\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=\mu(E_{t\bullet},\alpha_{\bullet},f_{t})=0$. In particular $E_{t\bullet}$ is $\Gamma$-invariant, and so is the associated filtration $V_{\bullet}$, which in turn yields $\mu(t_{3},\lambda)=0$. Hence, we have to show that $M(E_{t\bullet},\alpha_{\bullet})+\delta\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=0$ implies $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)=0$. This is the same calculation as for the moduli space of principal $G$-bundles with no twisted equivariant structure. For example, use a similar calculation to the one in the proof of [20, Theorem 2.3.5.16]. ∎
@@ 7.1. Finite torsors and moduli problems
The construction of $R_{1}$ is similar to [20, Theorem 2.4.8.7]. Given an open subspace $U$ of $R_{2}$, we denote by $\mathcal{F}_{2}^{U}$ the functor from the site of schemes over $U$ to sets, sending a morphism $f:S\to U$ to $\eta(S)^{-1}(f^{*}u_{2}|_{U\times X})$.
@@
In order to lift the $H$-action, we use the same argument as in the proof of [20, Proposition 2.4.8.9]. The $H$-action on $R_{2}$ is given by a morphism

### v2

@@ 6.3. The moduli space is a projective variety
By the proof of [21, Proposition 2.3.5.17], $\sigma_{C^{*}}$ extends to a family of homomorphisms $\sigma_{C}:(E_{C})_{a,b,c}\to\pi_{X}^{*}\mathcal{O}_{X}$. As in such proof, let $\iota:(C\times X)\setminus(C\times Z)\hookrightarrow C\times X$ be the open embedding and let
@@
Finally, as in the proof of [21, Proposition 2.3.5.17], one can check that $(\widetilde{E}_{0},\widetilde{\sigma}_{0},\widetilde{f}_{0})$ is a $(\delta,\chi)$-semistable pseudo-equivariant pair. Since its image in $\mathbb{G}{(m)}^{ss}$ is equal to $\eta(0)$, it determines a extension $\eta^{*}:C\to\mathfrak{C}_{0}^{ss}$ lifting $\eta$, as required. ∎
@@ 6.4. S-equivalence vs polystability
We claim that $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)=0$ implies that $M(E_{t\bullet},\alpha_{\bullet})+\delta\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=0$. This is just the same calculation as for the moduli space of principal $G$-bundles with no twisted equivariant structure. For example, use a similar calculation to the one in the proof of [21, Theorem 2.3.5.13] after the statement of [21, Corollary 2.3.5.14].
@@
Now assume that $M(E_{t\bullet},\alpha_{\bullet})=\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=\mu(E_{t\bullet},\alpha_{\bullet},f_{t})=0$. In particular $E_{t\bullet}$ is $\Gamma$-invariant, and so is the associated filtration $V_{\bullet}$, which in turn yields $\mu(t_{3},\lambda)=0$. Hence, we have to show that $M(E_{t\bullet},\alpha_{\bullet})+\delta\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=0$ implies $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)=0$. This is the same calculation as for the moduli space of principal $G$-bundles with no twisted equivariant structure. For example, use a similar calculation to the one in the proof of [21, Theorem 2.3.5.16]. ∎
@@ 7.1. Finite torsors and moduli problems
The construction of $R_{1}$ is similar to [21, Theorem 2.4.8.7]. Given an open subspace $U$ of $R_{2}$, we denote by $\mathcal{F}_{2}^{U}$ the functor from the site of schemes over $U$ to sets, sending a morphism $f:S\to U$ to $\eta(S)^{-1}(f^{*}u_{2}|_{U\times X})$.
@@
In order to lift the $H$-action, we use the same argument as in the proof of [21, Proposition 2.4.8.9]. The $H$-action on $R_{2}$ is given by a morphism